Suppose that a 5.0£10¡5 M solution of a compound had a %T equal to 70.80 when measured in a
cuvette with a 1.0 cm path length. Find themolar extinction coefficient of this compound.
1
Expert's answer
2020-09-22T03:36:08-0400

The term molar extinction coefficient (ε) is a measure of how strongly a chemical species or substance absorbs light at a particular wavelength. It is an intrinsic property of chemical species that is dependent upon their chemical composition and structure. The SI units of ε are m2/mol, but in practice they are usually taken as M-1cm-1. The molar extinction coefficient is frequently used in spectroscopy to measure the concentration of a chemical in solution.

You can use the Beer-Lambert Law to calculate a chemical species’ ε:

A = εLc

Where:

  • A is the amount of light absorbed by the sample for a particular wavelength
  • ε is the molar extinction coefficient
  • L is the distance that the light travels through the solution
  • c is the concentration of the absorbing species per unit volume

Rearrange the Beer-Lambert equation in order to solve for the molar extinction coefficient:

ε = A/Lc

Use the molar extinction coefficient to determine the brightness of a fluorescent molecule, by using the following equation:


So , ε = A/Lc = 70.80 / 1 * ( 5* 10^-5 ) = 1.5 * 10^6 . Answer .